Consider the standard normal random variable (Z) and answer questions 12 to 14.

Question 12

The area to the right of Z = 0 is _______________.
(1) 1
(2) The same as the area to the left of Z = 0
(3) Greater than the area to the left of Z = 0
(4) Less than the area to the left of Z = 0
(5) 0

Question 13

P(Z >= -2.88) = ?
(1) 0.0020
(2) 0.0014
(3) 0.9986
(4) 0.9980
(5) None of the above

Question 14

P(-1.1 <= Z <= 1.9) = ?

(1) 0.7285
(2) 0.8126
(3) 0.8356
(4) 0.9713
(5) None of the above
1
Expert's answer
2020-02-28T10:57:28-0500

12. The area to the right of Z=0 is (2) the same as the area to the left of Z = 0.

This is because Z = 0 is the point of symmetry in a standard normal curve.

13. P(Z>=-2.88)

P(Z>=-2.88)= P(Z<=2.88)

From the standard normal tables the value is 0.9980

The answer is (4)

14. P(-1.1<=Z <=1.9)

P(Z<=1.9)-P(Z<=-1.1)

P(Z<=1.9)-(1-P(Z<=1.1))

Reading from the standard normal tables the value of 1.1 and 1.9

=0.97128-(1-0.86433)

= 0.8356

The answer is (3)